# Rethinking Fine-Tuning when Scaling Test-Time Compute: Limiting Confidence Improves Mathematical Reasoning

### v3

@@ 5.2 DCO can prevent overconfidence and rescue test-time scaling
We next test whether DCO can rescue test-time scaling by preventing model overconfidence. We first perform experiments on the MATH dataset in this section, and then in the next section we perform experiments on the LeanDojo automated theorem proving benchmark [42]. We fine-tune Llama-3-8B base model for 4 epochs on the MATH training set using DCO for $N^{\prime}=256$ and confirm that the model is much less confident in its greedy completion than when trained with CE loss after multiple epochs (compare Figure 1 (a) and (c)). Moreover, training with DCO at larger $N$ yields lower model greedy confidences at the end of training (Figure 1 (d)), consistent with Lemma 4.2.
We next assess pass@N test performance as a function of $N$ for models trained by minimizing $\mathcal{L}_{\text{DCO}}^{N^{\prime}}$ for different values of $N^{\prime}$ (Figure 2 (a)). For any given $N$, there is an optimal $N^{\prime}$ for the training loss $\mathcal{L}_{\text{DCO}}^{N^{\prime}}$ that maximizes pass@N test coverage, yielding a Pareto optimal performance frontier (black curve) that is achieved when $N^{\prime}$ is close to $N$. In particular the model trained with CE loss (equivalent to pass@1 maximization at training) performs poorly relative to the Pareto frontier at large N (red curve below black at large $N$). Conversely, models trained with DCO at large $N^{\prime}$ perform poorly relative to the Pareto frontier for small $N$ (green curves below black at small $N$).
@@ 5.3 Improved theorem proving via ensembled tree search through a modified step-wise DCO
In our experiments, we use LEAN4 [44] as the formal proof language. We start from Qwen2.5-Math-1.5B [26], and fine-tune it on the LeanDojo benchmark [42]. In contrast to solving math problems above where the model first autoregressively generates an entire answer $y$ according to $\hat{p}(y|x)$ and then the entire $y$ is verified for correctness, in theorem proving we must obtain correct tactics $y[i]$, as verified by the proof assistant, at every proof step $i$. A straightforward application of DCO fine-tuning in this setting then involves replacing the model confidence $\hat{p}(y|x)$ in  Equation 3 on a single answer $y$, with the model confidence on an entire successful, complete $k$-step proof $\hat{p}(y[0],...,y[k-1]|x[0])=\prod_{i=0}^{k-1}\hat{p}(y[i]|x[i])$. Because of the chain rule, the DCO gradient on a single proof has the same form as Equation 4 but with $F(N,\hat{p}(y|x))$ replaced with $F(N,\prod_{i=0}^{k-1}\hat{p}\left(y[i]|x[i]\right))$. We naively applied this DCO fine-tuning method with $N^{\prime}=4$k and evaluated the resulting model, as well as a base model trained with CE loss, on MiniF2F, using pass@4k. CE loss achieves a proof success rate of 37.4%, while fine-tuning with DCO achieves a 38.7% success rate. Thus, a naive application of DCO to theorem proving by matching the parameter $N^{\prime}$ in DCO to the parameter $N$ in pass@N achieves only a modest improvement.
In theorem proving, since every single intermediate proof step tactic $y[i]$ must be valid, it is natural to consider a step-wise generalization of DCO. For example, at each step $i$, the model chooses a tactic $y[i]$ with confidence $\hat{p}(y[i]|x[i])$. Our proposed step-wise DCO optimizes this single-step confidence according to Equation 4 as before, except now the step-wise confidence regularizer is given by $F(N_{\text{eff}},\hat{p}(y[i]|x[i]))$. Here one can think of $N_{\text{eff}}$ as a step-wise DCO hyperparameter that controls the exploration width at every proof step. In essence, the confidence regularizer prevents the confidence of any chosen tactic from becoming much higher than $1/N_{\text{eff}}$. Since the sum of the confidences over all tactics at each step must be $1$, this means that at test time, model exploration at each step corresponds to searching on a search tree in which each proof state $x[i]$ allows the exploration of approximately only $N_{\text{eff}}$ tactics. Thus using $N_{\text{eff}}$ in step-wise DCO during fine-tuning selects the approximate branching factor $N_{\text{eff}}$ of the proof search tree at test time (see  Figure 2 (c)).
@@
In particular, a valid (partial) proof of length $k$ will have probability of order $N_{\text{eff}}^{-k}$. Thus small $N_{\text{eff}}$ limits the exploration at test time to a tree with small branching factor, but allows longer proofs with larger numbers of steps $k$ to have higher sampling probability. This limited width search strategy should work well at test time using pass@N as long as two conditions hold: (1) a successful proof of length $k$ is present in the search tree of branching factor $N_{\text{eff}}$; and (2) the $N$ in pass@N at test time is large enough that this proof of probability $O(N_{\text{eff}}^{-k})$ is selected with high probability in $N$ passes (which starts to occur as soon as $N\gtrsim N_{\text{eff}}^{k}$). Conversely, larger $N_{\text{eff}}$ allows for more exploration at test time using a wider search tree of larger branching factor, but it makes finding longer proofs with large $k$ harder, since the approximate success condition for pass@N of $N>N_{\text{eff}}^{k}$ is harder to satisfy at fixed $N$ and large $N_{\text{eff}}$ and $k$. However, this wide exploration strategy can work well for short proofs with small $k$ if a successful short proof does not lie in the limited width search tree obtained at small $N_{\text{eff}}$, but does lie in a wider search tree at larger $N_{\text{eff}}$ (Figure 2, (c) right).
The mean and median lengths of the proofs in the training set are 4.2 and 2, respectively. Thus we explore fine-tuning with the step-wise DCO algorithm, DCOstep, with $N_{\text{eff}}$ ranging from $1$ (CE loss) to $N_{\text{eff}}=32$. We evaluate pass@4k performance on the Mathlib and MiniF2F test sets and find that larger $N_{\mathrm{eff}}$ improves accuracy over the CE baseline (Table 2). Interestingly, the optimal $N_{\text{eff}}$ increases with more passes at test time (Table 6), similar to the Pareto frontier in Figure 2 (a). Since different choices of $N_{\text{eff}}$ correspond to different search strategies at test time—larger $N_{\text{eff}}$ favors short proofs in wide trees, smaller $N_{\text{eff}}$ favors long proofs in narrow trees— we hypothesized that ensembling these methods could significantly boost performance. This was indeed the case (ensemble entry in Table 2). The ensemble strategy samples $5\times 4$k times, so a proper baseline is CE loss with the same test-time compute of pass@N with $N=20$k. The ensemble strategy outperforms this stringent baseline with significant excesses of $5.2\%$ on Mathlib and $4.1\%$ on MiniF2F.
@@ 5.4 Approximate DCO also improves performance with chain-of-thought reasoning
Figure 3: (a, b) Overconfidence persists in CoT fine-tuning with CE loss. We fine-tune a Llama-3-8B base model on CoT traces in the MATH dataset and plot the distribution of the estimated model confidences $\hat{p}(y^{\mathrm{mode}}|x)$ over samples in the test set at various points in training. (a) For CE loss, the model becomes more confident in its most likely answers as training progresses—the confidence distribution shifts to the right. This effect is milder than in the direct answer setting (Figure 1). (b) DCOa limits this shift of the confidence distribution over training. (c, d) Overconfidence is also present in GRPO fine-tuning and can result in a model that is overconfident and wrong. (c) The model trained with GRPO assigns progressively larger confidences to $\hat{p}(y^{\mathrm{mode}}|x)$ over the course of training. (d) At the end of training, only a small portion of the model’s highly confident completions are correct, which will harm the model’s pass@N performance when scaling up $N$.
@@
In Section 5.2, the model is trained to directly give an answer $y$ to a problem $x$. In this case, one can implement DCO at training time by explicitly computing the model confidence $\hat{p}(y|x)$ and using it in Equation 4. However, in a Chain-of-Thought (CoT) training paradigm, the training data consists of triplets $(x,c,y)$ where $x$ and $y$ are the problem and answer as before, but now $c$ is a CoT reasoning trace that explains how to derive $y$ from $x$. The model is trained on the triplet $(x,c,y)$, and at test time, given a new problem $x^{\prime}$, it generates a reasoning trace $c^{\prime}$ followed by an answer $y^{\prime}$. Importantly, the model is evaluated on whether its answer $y^{\prime}$ is correct independent of its reasoning trace $c^{\prime}$.
@@
Thus, to estimate the probability $\hat{p}(y|x)$ that the model assigns to any answer $y$, one can no longer compute it directly as in the direct answer case. Instead, one must marginalize over all possible reasoning traces $c$ to obtain $\hat{p}(y|x)=\sum_{c}\hat{p}(y,c|x)$. Exact marginalization is intractable, so we replace the marginalization with a Monte Carlo estimate of $\hat{p}(y|x)$, which we insert into the overconfidence regularization factor $F$ in Equation 4. We call this algorithm approximate DCO, denoted by DCOa. We also compute Monte Carlo estimates of the probability the model assigns to its most likely final answer, $\hat{p}(y^{\mathrm{mode}}|x)$, as a proxy for the model’s max confidence. Note that in the CoT setting the greedy answer would correspond to a single most likely reasoning trace, but we want the probability of the most likely final answer marginalizing over all reasoning traces.
@@
We conduct experiments on the MATH dataset. As a baseline, we first fine-tune a Llama-3-8B-base model using CE loss on the golden CoT solutions. We find in Table 3 that pass@1 coverage robustly improves over the course of CE loss training. Intriguingly, in the CoT setting, overfitting induced by the misalignment of CE loss at training time with the pass@N strategy at test time, appears less severe compare with the direct answer setting (cf. Table 1). Nevertheless, we still find a clear increase in the model’s max confidence through training (Figure 3 (a)). In contrast, training with DCOa mitigates the rise in the model’s max confidence (Figure 3 (b)). As with the direct answer setup, model trained with DCOa ($N^{\prime}=64$) exhibits lower pass@1 performance compared to the CE baseline but demonstrates superior performance when evaluated with pass@N for larger $N$ values (Table 3).
@@
To provide a comprehensive comparison, we also evaluate a model trained with reinforcement learning. Specifically, we use GRPO [27] to train the base model with the r1 [10] prompt template. Interestingly, we observe that the GRPO-trained model exhibits pronounced overconfidence (Figure 3 (c)), yet only a small portion of the model’s highly confident answers are correct (Figure 3 (d)). This negatively impacts the model’s pass@N performance when scaling up $N$ (Table 3). Although GRPO demonstrates superior pass@1 performance, the model trained with DCOa achieves a significantly stronger performance when evaluated with pass@N for larger $N$ values (Table 3).
@@
Table 3: Pass@N coverage on MATH ($\pm$ indicates standard error of the mean). We fine-tune Llama-3-8B-base on CoT traces. Training with DCOa using $N^{\prime}=64$ underperforms CE loss for pass@1, but outperforms it for higher $N$, with the largest improvement when $N^{\prime}=N$. The GRPO model has the strongest pass@1 performance, but its performance degrades at larger $N$ due to overconfidence.
@@ 6 Discussion
We would like to thank Zhinan Cheng, Clémentine Dominé, Marco Fumero, David Klindt, Daniel Kunin, Ben Sorscher and Atsushi Yamamura for helpful discussions. S.G. thanks the James S. McDonnell Foundation, Simons Foundation, NTT Research, Schmidt Foundation and an NSF CAREER Award for support. S.D. and F.C. were partially supported by the McKnight Foundation, the Simons Foundation and an NIH CRCNS grant R01DC020874.
@@ A.1 Proof of approximately well calibration under optimal policy
Given any problem, let $\hat{p}_{i}$ denote the model probability or confidence assigned to answer $i$, and assume answers are sorted from highest to lowest confidence so that $\hat{p}_{i}\geq\hat{p}_{i+1},\forall\,i\geq 1$. Let $p_{i}$ be the probability that answer $i$ is actually correct across all the problems. Then optimal policies maximizing pass@N coverage in Equation 2 (denoted by $\hat{p}_{i}^{*}$) is approximately well calibrated, i. e. higher confidence implies higher or equal probability of being correct, i.e. $p_{i}\geq p_{i+1},\forall\,i\geq 1$.
@@
Let $R$ be the total number of answers the model can generate. Here $R$ could possibly equal infinity. Let $\hat{p}_{i}^{*}$ denote the model probability or confidence assigned to answer $i$ under optimal policy maximizing pass@N coverage in Equation 2. Let $A_{i}$ be the event that the $i$th ranked answer is not chosen in the $N$ passes and $B_{i}$ be the event that $i$th ranked answer is chosen in the $N$ passes. Let $1_{B_{i}}$ be the indicator function of $B_{i}$. Then, given an event $\omega$, the probability of getting the correct answer is $\sum_{i=1}^{R}p_{i}1_{B_{i}}(\omega)$. Hence, the expected probability of getting the correct answer is

### v4

@@ 5.2 DCO can prevent overconfidence and rescue test-time scaling
We next test whether DCO can rescue test-time scaling by preventing model overconfidence. We first perform experiments on the MATH dataset in this section, and then in the next section we perform experiments on the LeanDojo automated theorem proving benchmark [44]. We fine-tune Llama-3-8B base model for 4 epochs on the MATH training set using DCO for $N^{\prime}=256$ and confirm that the model is much less confident in its greedy completion than when trained with CE loss after multiple epochs (compare Figure˜1 (a) and (c)). Moreover, training with DCO at larger $N$ yields lower model greedy confidences at the end of training (Figure˜1 (d)), consistent with Lemma˜4.2.
We next assess pass@N test performance as a function of $N$ for models trained by minimizing $\mathcal{L}_{\text{DCO}}^{N^{\prime}}$ for different values of $N^{\prime}$ (Figure˜2 (a)). For any given $N$, there is an optimal $N^{\prime}$ for the training loss $\mathcal{L}_{\text{DCO}}^{N^{\prime}}$ that maximizes pass@N test coverage, yielding a Pareto optimal performance frontier (black curve) that is achieved when $N^{\prime}$ is close to $N$. In particular the model trained with CE loss (equivalent to pass@1 maximization at training) performs poorly relative to the Pareto frontier at large N (red curve below black at large $N$). Conversely, models trained with DCO at large $N^{\prime}$ perform poorly relative to the Pareto frontier for small $N$ (green curves below black at small $N$).
@@ 5.3 Improved theorem proving via ensembled tree search through a modified step-wise DCO
In our experiments, we use LEAN4 [46] as the formal proof language. We start from Qwen2.5-Math-1.5B [28], and fine-tune it on the LeanDojo benchmark [44]. In contrast to solving math problems above where the model first autoregressively generates an entire answer $y$ according to $\hat{p}(y|x)$ and then the entire $y$ is verified for correctness, in theorem proving we must obtain correct tactics $y[i]$, as verified by the proof assistant, at every proof step $i$. A straightforward application of DCO fine-tuning in this setting then involves replacing the model confidence $\hat{p}(y|x)$ in  Equation˜3 on a single answer $y$, with the model confidence on an entire successful, complete $k$-step proof $\hat{p}(y[0],...,y[k-1]|x[0])=\prod_{i=0}^{k-1}\hat{p}(y[i]|x[i])$. Because of the chain rule, the DCO gradient on a single proof has the same form as Equation˜4 but with $F(N,\hat{p}(y|x))$ replaced with $F(N,\prod_{i=0}^{k-1}\hat{p}\left(y[i]|x[i]\right))$. We naively applied this DCO fine-tuning method with $N^{\prime}=4$k and evaluated the resulting model, as well as a base model trained with CE loss, on MiniF2F, using pass@4k. CE loss achieves a proof success rate of 37.4%, while fine-tuning with DCO achieves a 38.7% success rate. Thus, a naive application of DCO to theorem proving by matching the parameter $N^{\prime}$ in DCO to the parameter $N$ in pass@N achieves only a modest improvement.
In theorem proving, since every single intermediate proof step tactic $y[i]$ must be valid, it is natural to consider a step-wise generalization of DCO. For example, at each step $i$, the model chooses a tactic $y[i]$ with confidence $\hat{p}(y[i]|x[i])$. Our proposed step-wise DCO optimizes this single-step confidence according to Equation˜4 as before, except now the step-wise confidence regularizer is given by $F(N_{\text{eff}},\hat{p}(y[i]|x[i]))$. Here one can think of $N_{\text{eff}}$ as a step-wise DCO hyperparameter that controls the exploration width at every proof step. In essence, the confidence regularizer prevents the confidence of any chosen tactic from becoming much higher than $1/N_{\text{eff}}$. Since the sum of the confidences over all tactics at each step must be $1$, this means that at test time, model exploration at each step corresponds to searching on a search tree in which each proof state $x[i]$ allows the exploration of approximately only $N_{\text{eff}}$ tactics. Thus using $N_{\text{eff}}$ in step-wise DCO during fine-tuning selects the approximate branching factor $N_{\text{eff}}$ of the proof search tree at test time (see  Figure˜2 (c)).
@@
In particular, a valid (partial) proof of length $k$ will have probability of order $N_{\text{eff}}^{-k}$. Thus small $N_{\text{eff}}$ limits the exploration at test time to a tree with small branching factor, but allows longer proofs with larger numbers of steps $k$ to have higher sampling probability. This limited width search strategy should work well at test time using pass@N as long as two conditions hold: (1) a successful proof of length $k$ is present in the search tree of branching factor $N_{\text{eff}}$; and (2) the $N$ in pass@N at test time is large enough that this proof of probability $O(N_{\text{eff}}^{-k})$ is selected with high probability in $N$ passes (which starts to occur as soon as $N\gtrsim N_{\text{eff}}^{k}$). Conversely, larger $N_{\text{eff}}$ allows for more exploration at test time using a wider search tree of larger branching factor, but it makes finding longer proofs with large $k$ harder, since the approximate success condition for pass@N of $N>N_{\text{eff}}^{k}$ is harder to satisfy at fixed $N$ and large $N_{\text{eff}}$ and $k$. However, this wide exploration strategy can work well for short proofs with small $k$ if a successful short proof does not lie in the limited width search tree obtained at small $N_{\text{eff}}$, but does lie in a wider search tree at larger $N_{\text{eff}}$ (Figure˜2, (c) right).
The mean and median lengths of the proofs in the training set are 4.2 and 2, respectively. Thus we explore fine-tuning with the step-wise DCO algorithm, DCOstep, with $N_{\text{eff}}$ ranging from $1$ (CE loss) to $N_{\text{eff}}=32$. We evaluate pass@4k performance on the Mathlib and MiniF2F test sets and find that larger $N_{\mathrm{eff}}$ improves accuracy over the CE baseline (Table˜2). Interestingly, the optimal $N_{\text{eff}}$ increases with more passes at test time (Table˜6), similar to the Pareto frontier in Figure˜2 (a). Since different choices of $N_{\text{eff}}$ correspond to different search strategies at test time—larger $N_{\text{eff}}$ favors short proofs in wide trees, smaller $N_{\text{eff}}$ favors long proofs in narrow trees— we hypothesized that ensembling these methods could significantly boost performance. This was indeed the case (ensemble entry in Table˜2). The ensemble strategy samples $5\times 4$k times, so a proper baseline is CE loss with the same test-time compute of pass@N with N=20N=20k. The ensemble strategy outperforms this stringent baseline with significant excesses of $5.2\%$ on Mathlib and $4.1\%$ on MiniF2F.
@@ 5.4 Approximate DCO also improves performance with chain-of-thought reasoning
Figure 3: (a, b) Overconfidence persists in CoT fine-tuning with CE loss. We fine-tune a Llama-3-8B base model on CoT traces in the MATH dataset and plot the distribution of the estimated model confidences $\hat{p}(y^{\mathrm{mode}}|x)$ over samples in the test set at various points in training. (a) For CE loss, the model becomes more confident in its most likely answers as training progresses—the confidence distribution shifts to the right. This effect is milder than in the direct answer setting (Figure˜1). (b) DCOa limits this shift of the confidence distribution over training. (c, d) Overconfidence is also present in GRPO fine-tuning and can result in a model that is overconfident and wrong. (c) The model trained with GRPO assigns progressively larger confidences to $\hat{p}(y^{\mathrm{mode}}|x)$ over the course of training. (d) At the end of training, only a small portion of the model’s highly confident completions are correct, which will harm the model’s pass@N performance when scaling up NN.
@@
In Section˜5.2, the model is trained to directly give an answer $y$ to a problem $x$. In this case, one can implement DCO at training time by explicitly computing the model confidence $\hat{p}(y|x)$ and using it in Equation˜4. However, in a Chain-of-Thought (CoT) training paradigm, the training data consists of triplets $(x,c,y)$ where $x$ and $y$ are the problem and answer as before, but now $c$ is a CoT reasoning trace that explains how to derive $y$ from $x$. The model is trained on the triplet $(x,c,y)$, and at test time, given a new problem $x^{\prime}$, it generates a reasoning trace $c^{\prime}$ followed by an answer $y^{\prime}$. Importantly, the model is evaluated on whether its answer $y^{\prime}$ is correct independent of its reasoning trace $c^{\prime}$.
@@
Thus, to estimate the probability $\hat{p}(y|x)$ that the model assigns to any answer $y$, one can no longer compute it directly as in the direct answer case. Instead, one must marginalize over all possible reasoning traces $c$ to obtain $\hat{p}(y|x)=\sum_{c}\hat{p}(y,c|x)$. Exact marginalization is intractable, so we replace the marginalization with a Monte Carlo estimate of $\hat{p}(y|x)$, which we insert into the overconfidence regularization factor $F$ in Equation˜4. We call this algorithm approximate DCO, denoted by DCOa. We also compute Monte Carlo estimates of the probability the model assigns to its most likely final answer, $\hat{p}(y^{\mathrm{mode}}|x)$, as a proxy for the model’s max confidence. Note that in the CoT setting the greedy answer would correspond to a single most likely reasoning trace, but we want the probability of the most likely final answer marginalizing over all reasoning traces.
@@
We conduct experiments on the MATH dataset. As a baseline, we first fine-tune a Llama-3-8B-base model using CE loss on the golden CoT solutions. We find in Table˜3 that pass@1 coverage robustly improves over the course of CE loss training. Intriguingly, in the CoT setting, overfitting induced by the misalignment of CE loss at training time with the pass@N strategy at test time, appears less severe compare with the direct answer setting (cf. Table˜1). Nevertheless, we still find a clear increase in the model’s max confidence through training (Figure˜3 (a)). In contrast, training with DCOa mitigates the rise in the model’s max confidence (Figure˜3 (b)). As with the direct answer setup, model trained with DCOa ($N^{\prime}=64$) exhibits lower pass@1 performance compared to the CE baseline but demonstrates superior performance when evaluated with pass@N for larger $N$ values (Table˜3).
@@
To provide a comprehensive comparison, we also evaluate a model trained with reinforcement learning. Specifically, we use GRPO [29] to train the base model with the r1 [10] prompt template. Interestingly, we observe that the GRPO-trained model exhibits pronounced overconfidence (Figure˜3 (c)), yet only a small portion of the model’s highly confident answers are correct (Figure˜3 (d)). This negatively impacts the model’s pass@N performance when scaling up $N$ (Table˜3). Although GRPO demonstrates superior pass@1 performance, the model trained with DCOa achieves a significantly stronger performance when evaluated with pass@N for larger $N$ values (Table˜3).
@@
Table 3: Pass@N coverage on MATH ($\pm$ indicates standard error of the mean). We fine-tune Llama-3-8B-base on CoT traces. Training with DCOa using $N^{\prime}=64$ underperforms CE loss for pass@1, but outperforms it for higher NN, with the largest improvement when $N^{\prime}=N$. The GRPO model has the strongest pass@1 performance, but its performance degrades at larger NN due to overconfidence.
@@ 6 Discussion
We would like to thank Zhinan Cheng, Clémentine Dominé, Marco Fumero, David Klindt, Daniel Kunin, Ben Sorscher and Atsushi Yamamura for helpful discussions. S.G. thanks the James S. McDonnell Foundation, Simons Foundation, NTT Research, Schmidt Foundation and an NSF CAREER Award for support. S.D. and F.C. were partially supported by the McKnight Foundation, the Simons Foundation and an NIH CRCNS grant R01DC020874. We are also grateful to Adam Ousherovitch and Hongda Yuan for identifying an error in Lemma˜4.2, which has been corrected in the revised version of this paper.
@@ A.1 Proof of approximately well calibration under optimal policy
Given any problem, let $\hat{p}_{i}$ denote the model probability or confidence assigned to answer $i$, and assume answers are sorted from highest to lowest confidence so that $\hat{p}_{i}\geq\hat{p}_{i+1},\forall\,i\geq 1$. Let $p_{i}$ be the probability that answer $i$ is actually correct across all the problems. Then optimal policies maximizing pass@N coverage in Equation˜2 (denoted by $\hat{p}_{i}^{*}$) is approximately well calibrated, i. e. higher confidence implies higher or equal probability of being correct, i.e. $p_{i}\geq p_{i+1},\forall\,i\geq 1$.
@@
Let $R$ be the total number of answers the model can generate. Here $R$ could possibly equal infinity. Let $\hat{p}_{i}^{*}$ denote the model probability or confidence assigned to answer $i$ under optimal policy maximizing pass@N coverage in Equation˜2. Let $A_{i}$ be the event that the $i$th ranked answer is not chosen in the $N$ passes and $B_{i}$ be the event that $i$th ranked answer is chosen in the $N$ passes. Let $1_{B_{i}}$ be the indicator function of $B_{i}$. Then, given an event $\omega$, the probability of getting the correct answer is $\sum_{i=1}^{R}p_{i}1_{B_{i}}(\omega)$. Hence, the expected probability of getting the correct answer is
@@
Assuming $p_{1}$ is known, an upper bound on $\hat{p}_{1}^{*}$ is given by,
$u=\underset{\{p_{s}\}_{s=1}^{k}}{\max}\pi_{1}\left(\underset{\{\hat{p}_{i}\}_{i=1}^{k}}{\text{argmin}}\left\{\sum_{i=1}^{k}p_{i}(1-\hat{p}_{i})^{N}\right\}\right)$ (8)
$\text{s.t.}\;\sum_{s=1}^{k}p_{s}=1-\epsilon,\;p_{1}\geq p_{2}.\geq p_{k}>0.$
Suppose $\{p_{i}\}_{i=1}^{k}$ solves Equation˜8, and let $\{\hat{p}_{i}^{*}\}_{i=1}^{k}$ be the optimizer of
$\underset{\{\hat{p}_{i}\}_{i=1}^{k}}{\text{argmin}}\{\sum_{i=1}^{k}p_{i}(1-\hat{p}_{i})^{N}\}\;\text{s.t.}\;\sum_{s=1}^{k}\hat{p}_{s}=1,\;1\geq\hat{p}_{1}\geq 0,...,1\geq\hat{p}_{k}\geq 0.$ (9)
By definition, $u=\hat{p}_{1}^{*}$. Since $p_{1}\geq p_{2}...\geq p_{k}$, we also have $\hat{p}_{1}^{*}\geq\hat{p}_{2}^{*}...\geq\hat{p}_{k}^{*}$. There are $k$ possible cases.
  • 1)
    $\hat{p}_{1}^{*}=1$ and $\hat{p}_{s}^{*}=0$ for all $s\geq 2$.
  • $j$)
    $\hat{p}_{j}^{*}>0$ but $\hat{p}_{j+1}^{*}=0$. For $2\leq j\leq k-1$. There are k2k-2 cases.
  • $k$)
    $\hat{p}_{k}^{*}>0$.
We now characterize when each case can occur and derive an upper bound in each case.
A.2.1 Case $k$
Lemma A.2.
If $\hat{p}_{k}^{*}>0$, then $p_{1}=p_{2}=...=p_{k-1}$.